OpenStudy (anonymous):

1/x^2+4x+3+ 1/ x^2-1 ...please help me solve this

OpenStudy (anonymous):

is it this\[\huge \frac{1}{x^2} +4x + 3 + \frac{1}{x^2-1}\]?

OpenStudy (anonymous):

no the 4x+3 goes under 1 in the first fraction

OpenStudy (anonymous):

\[\frac{1}{x^2+4x+3}+ \frac{1}{x^2-1}\]right?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

let's see what i can do with this, just simplify i guess first factor \[x^2 + 4x + 3 \space and \space x^2 - 1\]

OpenStudy (anonymous):

(x+3) (x+1) and (x+1) (x-1)

OpenStudy (anonymous):

\[\frac{1}{(x+3)(x+1)} + \frac{1}{(x+1)(x-1)}\]make the denominators common~

OpenStudy (anonymous):

that's what I don't get..I'm stuck there

OpenStudy (anonymous):

\[\frac{(x-1)+(x+3)}{(x+3)(x+1)(x-1)}\]rings any bells? :s

OpenStudy (anonymous):

sort of

OpenStudy (anonymous):